Frequency Table showing distribution of Exam scores
Exam scores # of students
90 & over 5
80-89 12
70-79 40
60-69 18
50-59 13
40-49 6
30-39 6
What is the probability that a student picked at random will have a scor of at least 90? a score less than 90? a score of at least 60? and a score of more than 39 and less than 60? Tried several different equations with this and it just looked crazy;P(<90)= .12 is this correct?

Answer by funmath(2925)

Frequency Table showing distribution of Exam scores
Exam scores # of students
90 & over 5
80-89 12
70-79 40
60-69 18
50-59 13
40-49 6
30-39 6
What is the probability that a student picked at random will have a score of at least 90? a score less than 90? a score of at least 60? and a score of more than 39 and less than 60?
There are a total of 5+12+40+18+13+6+6=100 students
P of at least 90 is 5/100=.05
P of less than 90 is (12+40+18+13+6+6)/100=.95
P of at least 60 is (5+12+40+18)/100=.75
P of more than 39 and less than 60 is (13+6+)/100=.19
:
Happy Calculating!!!